OpenStudy (anonymous):

write 0.35 as fraction or mixed number

OpenStudy (solomonzelman):

\[.35=\frac{35}{100}=\frac{7}{20}\]

OpenStudy (charlotte123):

You're good man ლ(́◉◞౪◟◉‵ლ)

OpenStudy (solomonzelman):

(it's a fraction, mixed number is equal to more than 1)

OpenStudy (charlotte123):

Thanks for the medal! :D

OpenStudy (anonymous):

Write the decimal number as a fraction (over 1) 0.35 = 0.35 / 1 Multiplying by 1 to eliminate 2 decimal places we multiply top and bottom by 2 10's Numerator (N) N = 0.35 × 10 × 10 = 35 Denominator (D) D = 1 × 10 × 10 = 100 N / D = 35 / 100 Simplifying our fraction = 35/100 = 7/20

OpenStudy (anonymous):

is mine better

OpenStudy (anonymous):

have anymore of those

OpenStudy (charlotte123):

Great explanation Angel!

OpenStudy (anonymous):

welcome and thanks!

OpenStudy (anonymous):

anymore

OpenStudy (anonymous):

@kimmycute do you have anymore of those

OpenStudy (anonymous):

ok then bye

OpenStudy (anonymous):

i am the best one at those you know i do it step by step

OpenStudy (anonymous):

does she need help anymore?

OpenStudy (anonymous):

don't know

OpenStudy (anonymous):

ok.do you need help ;)

OpenStudy (anonymous):

@angel12310

OpenStudy (charlotte123):

LOLLLLLLLLLLLLL

OpenStudy (charlotte123):

My answer disappeared...........weird O_O

OpenStudy (anonymous):

what so funny @charlotte123

OpenStudy (charlotte123):

Do you need help *insert wink face here* Dude......THAT WAS HILARIOUS xD

OpenStudy (anonymous):

@angel12310 so wat up wit you?

OpenStudy (anonymous):

i don't need help

OpenStudy (anonymous):

@kimmycute